# Nash Equilibrium between Brokers and Traders

### v1

@@ Abstract
We study the perfect information Nash equilibrium between a broker and her clients — an informed trader, and an uniformed trader. In our model, the broker trades in the lit exchange where trades have instantaneous and transient price impact with exponential resilience, while both clients trade with the broker. The informed trader and the broker maximise expected wealth subject to inventory penalties, while the uninformed trader is not strategic and sends the broker random buy and sell orders. We characterise the Nash equilibrium of the trading strategies with the solution to a coupled system of forward-backward stochastic differential equations (FBSDEs). We solve this system explicitly and study the effect of information in the trading strategies of the broker and the informed trader.
@@ 1 Introduction
The remainder of the paper proceeds as follows. Section 2 introduces the model, the performance criteria, and defines the notion of a Nash equilibrium in our setup. Section 3 characterises the Nash equilibrium of the game and solves it explicitly. First, we show that both functionals are strictly concave (up to negligible sets), then, we compute their Gâteaux derivatives, we characterise the Nash equilibrium with a system of FBSDEs, and we study existence and uniqueness of the system. Section 4 derives the closed-form solution and connects the existence and uniqueness of the FBSDE with the existence and uniqueness of a matrix Riccati differential equation; we show existence and uniqueness of this differential equation under some restrictions of model parameters. Finally, Section 5 shows numerical experiments and concludes.
@@ 2 The model
Sometimes we write the control in the superscript when we highlight the controls that affect a process, for example, we write $X^{I,\eta}$ and $Q^{I,\eta}$ for the cash process and inventory process of the informed trader. The sets of admissible strategies for the informed trader and the broker are both given by the set ${\mathbb{H}^{2}}$ of squared-integrable adapted processes.
@@ 3.1 Characterisation of equilibrium
First, we show that for fixed $\nu\in{\mathbb{H}^{2}}$ the functional $\eta\to J^{I}(\nu,\eta)$ is strictly concave up to negligible sets; this is proven in the next proposition.
@@
$2\,b\,\eta^{*}_{t}$ $=-2\,\psi\,Q^{I,\eta^{*}}_{t}+\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}$ (3.30)
up to $\mathbb{P}\otimes\mathbb{L}$ null sets. Next, define the ${\mathcal{F}}$-martingale $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ by
@@
Sufficiency: Let $\eta^{*}\in{\mathbb{H}^{2}}$ satisfy (LABEL:eq:_FBSDE_1_Informed). The unique strong solution to this linear FBSDE is
@@
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}-M^{I}_{t}+M^{I}_{T}\bigg{\}}\,\mathrm{d}t\Bigg{]}=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}\mathbb{E}\bigg{[}-M^{I}_{t}+M^{I}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\bigg{]}\bigg{\}}\,\mathrm{d}t\Bigg{]}=0\,,$
@@
$2\,a\,\nu^{*}_{t}$ $=-(2\,\phi-\mathfrak{h})\,Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\,\mathbb{E}\left[\int_{t}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]$
up to $\mathbb{P}\otimes\mathbb{L}$ null sets. Hence, define the ${\mathcal{F}}$-martingales $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$ s.t.,
@@
and together with (3.1) evaluated at $T$ show that the FBSDE system is satisfied. Sufficiency: Let $\nu^{*}\in{\mathbb{H}^{2}}$ satisfy (3.38). The unique strong solution to this linear FBSDE is
@@
where $\varphi=\phi-\tfrac{1}{2}\mathfrak{h}$, and the processes $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$ are ${\mathcal{F}}$-martingales, and $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ is a ${\mathcal{F}}$-martingale. Then, $(\nu^{*},\eta^{*})\in{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}$ is a Nash equilibrium of the trading strategies between the broker and the informed trader.
@@
Thus, $(\eta^{*},\nu^{*})$ are a Nash equilibrium because they satisfy Definition 2.2. Uniqueness up to $\mathbb{P}\otimes\mathbb{L}$ null sets follows from the strict concavity of both $\eta\to J^{I}(\eta,\nu^{*})$ and $\nu\to J^{B}(\eta^{*},\nu)$. ∎
@@ 3.2 Existence and uniqueness
The FBSDE in (3.47) can be written compactly as follows
@@
If $T$ is small enough, then, existence and uniqueness of the slution to the FBSDE (3.51) is well understood in the literature; see e.g., Antonelli, (1993) or Chapter 2 in Carmona, (2016). In the Brownian case, Yong, (1999) shows that existence and uniqueness of a solution to the FBSDE system (3.51) is equivalent to the existence and uniqueness of a matrix Riccati differential equation; this result follows the ‘four-step scheme’ of Ma et al., (1994).222 See Ma and Yong, (1999); Carmona, (2016) for manuscripts covering a wide range of existence and uniqueness results for FBSDEs. Below, similar to these previous works, we take a constructive approach and show that existence and uniqueness of the solution to the FBSDE (3.51) is equivalent to the existence and uniqueness of a non-Hermitian matrix Riccati differential equation. We show that when the values of the parameters $\mathfrak{p}$ and $r^{B}$ are zero, existence and uniqueness of the solution follows without any further conditions on the remaining model parameters. Alternatively, if only $\mathfrak{p}$ is zero, existence and uniqueness follows under four additional (mild) restrictions on model parameters. In the numerical examples below we take $\mathfrak{p}=r^{B}=0$, thus, we employ the result of existence and uniqueness that we prove below with tools from the literature on non-Hermitian matrix Riccati differential equations. Finally, in the general case when $\mathfrak{p}$ is not zero and $T$ is small enough, then existence and uniqueness follow from the arguments in Antonelli, (1993).
@@ 5 Numerical results
The mean mark-to-market (MtM) of the informed trader increases by roughly 4.5% when going from the two-stage optimisation in Cartea and Sánchez-Betancourt, (2022) to the Nash equilibrium, whereas the broker decreases their mean MtM by 5.5%. A $t$-test shows, with 99% confidence, that the average MtM of the samples generated by both the Nash and two-stage optimisation are statistically different (for both the informed trader and the broker). This difference provides an upper bound for how profitability moves between players when information becomes publicly available. Thus, in our model (for the parameters above), the maximum improvement that the informed trader can obtain is roughly 5% which is largely paid by the broker; note that this is four orders of magnitude higher than transaction costs in foreign exchange markets. [^5] Transaction costs in spot foreign exchange are roughly \$3/\$ M; see Cartea and Sánchez-Betancourt, (2023). Our results provide insights into what happens in broker-client relationships in markets where the broker does not benefit from anonymity. Such perfect information trading environments are a reality in the decentralised finance space – in particular in automated market maker trading venues.

### v2

@@ Abstract
We study the perfect information Nash equilibrium between a broker and her clients — an informed trader and an uniformed trader. In our model, the broker trades in the lit exchange where trades have instantaneous and transient price impact with exponential resilience, while both clients trade with the broker. The informed trader and the broker maximise expected wealth subject to inventory penalties, while the uninformed trader is not strategic and sends the broker random buy and sell orders. We characterise the Nash equilibrium of the trading strategies with the solution to a coupled system of forward-backward stochastic differential equations (FBSDEs). We solve this system explicitly and study the effect of information, profitability, and inventory control in the trading strategies of the broker and the informed trader.
@@ 1 Introduction
The remainder of the paper proceeds as follows. Section 2 introduces the model, the performance criteria, and defines the notion of a Nash equilibrium in our setup. Section 3 characterises the Nash equilibrium of the game and solves it explicitly. First, we show that both functionals are strictly concave, then, we compute their Gâteaux derivatives, we characterise the Nash equilibrium with a system of FBSDEs, and we study existence and uniqueness of the system. Section 4 derives the closed-form solution and connects the existence and uniqueness of the FBSDE with the existence and uniqueness of a matrix Riccati differential equation; we show existence and uniqueness of this differential equation under some restrictions of model parameters. Finally, Section 5 shows numerical experiments and concludes.
@@ 2 The model
Sometimes we write the control in the superscript when we highlight the controls that affect a process, for example, we write $X^{I,\eta}$ and $Q^{I,\eta}$ for the cash process and inventory process of the informed trader. The sets of admissible strategies for the informed trader and the broker are both given by the set ${\mathbb{H}^{2}}$.
@@ 3.1 Characterisation of equilibrium
First, we show that for fixed $\nu\in{\mathbb{H}^{2}}$ the functional $\eta\to J^{I}(\nu,\eta)$ is strictly concave; this is proven in the next proposition.
@@
$2\,b\,\eta^{*}_{t}$ $=-2\,\psi\,Q^{I,\eta^{*}}_{t}+\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}\,.$ (3.30)
Next, define the ${\mathcal{F}}$-martingale $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ by
@@
Sufficiency: Let $\eta^{*}\in{\mathbb{H}^{2}}$ satisfy (LABEL:eq:_FBSDE_1_Informed). The solution to the FBSDE can be written as
@@
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}-M^{I}_{t}+M^{I}_{T}\bigg{\}}\,\mathrm{d}t\Bigg{]}=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}\mathbb{E}\bigg{[}-M^{I}_{t}+M^{I}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\bigg{]}\bigg{\}}\,\mathrm{d}t\Bigg{]}=0\,.$
@@
$2\,a\,\nu^{*}_{t}$ $=-(2\,\phi-\mathfrak{h})\,Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\,\mathbb{E}\left[\int_{t}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]\,.$
Hence, define the ${\mathcal{F}}$-martingales $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$ s.t.,
@@
and together with (3.1) evaluated at $T$ show that the FBSDE system is satisfied. Sufficiency: Let $\nu^{*}\in{\mathbb{H}^{2}}$ satisfy (3.38). The solution to the FBSDE can be written as
@@
where $\varphi=\phi-\tfrac{1}{2}\mathfrak{h}$, and the processes $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$, $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$, and $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ are ${\mathcal{F}}$-martingales. Then, $(\nu^{*},\eta^{*})\in{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}$ is a Nash equilibrium of the trading strategies between the broker and the informed trader.
@@
Thus, $(\eta^{*},\nu^{*})$ are a Nash equilibrium because they satisfy Definition 2.2. Uniqueness follows from the strict concavity of both $\eta\to J^{I}(\eta,\nu^{*})$ and $\nu\to J^{B}(\eta^{*},\nu)$. ∎
@@ 3.2 Existence and uniqueness
The FBSDE in (3.47) can be written as follows
@@
If $T$ is small enough, then, existence and uniqueness of the solution to the FBSDE (3.51) is well understood in the literature; see e.g., Antonelli, (1993) or Chapter 2 in Carmona, (2016). In the Brownian case, Yong, (1999) shows that existence and uniqueness of a solution to the FBSDE system (3.51) is equivalent to the existence and uniqueness of a matrix Riccati differential equation; this result follows the ‘four-step scheme’ of Ma et al., (1994).222 See Ma and Yong, (1999); Carmona, (2016) for manuscripts covering a wide range of existence and uniqueness results for FBSDEs. Below, similar to these previous works, we take a constructive approach and show that existence and uniqueness of the solution to the FBSDE (3.51) is equivalent to the existence and uniqueness of a non-Hermitian matrix Riccati differential equation. We show that when the values of the parameters $\mathfrak{p}$ and $r^{B}$ are zero, existence and uniqueness of the solution follows without any further conditions on the remaining model parameters. Alternatively, if only $\mathfrak{p}$ is zero, existence and uniqueness follows under four additional (mild) restrictions on model parameters. In the numerical examples below we take $\mathfrak{p}=r^{B}=0$, thus, we employ the result of existence and uniqueness that we prove below with tools from the literature on non-Hermitian matrix Riccati differential equations. Finally, in the general case when $\mathfrak{p}$ is not zero and $T$ is small enough, then existence and uniqueness follow from the arguments in Antonelli, (1993).
@@ 5 Numerical results
The mean mark-to-market (MtM) of the informed trader increases by roughly 4.5% when going from the two-stage optimisation in Cartea and Sánchez-Betancourt, (2022) to the Nash equilibrium, whereas the broker decreases their mean MtM by 5.5%. A $t$-test shows, with 99% confidence, that the average MtM of the samples generated by both the Nash and two-stage optimisation are statistically different (for both the informed trader and the broker). This difference provides an upper bound for how profitability moves between players when information becomes publicly available. Thus, in our model (for the parameters above), the maximum improvement that the informed trader can obtain is roughly 5% which is largely paid by the broker; note that this is four orders of magnitude higher than transaction costs in foreign exchange markets. [^5] Transaction costs in spot foreign exchange are roughly \$3 per \$1 million traded; see Cartea and Sánchez-Betancourt, (2023). Our results provide insights into what happens in broker-client relationships in markets where the broker does not benefit from anonymity. Such perfect information trading environments are a reality in the decentralised finance space – in particular in automated market maker trading venues.